OpenStudy (gorica):

What is position vector equation of a simple pendulum using normal and tangent vector? String has length s and the angle that it makes with vertical axes is θ.

OpenStudy (anonymous):

i think we are to work with polar coordinates with origin at the point of suspension.

OpenStudy (gorica):

I would write what I have got, but I don't know how to use LaTex here. I can attach a picture to see what I wrote.

OpenStudy (anonymous):

i am having problem using latex too. I want to use unit vectors along length and along increasing value to \(\theta\). -these are unit vectors in polar coordinates.But i don't know how to put "cap" on them!

OpenStudy (gorica):

Do you agree to write ' instead of cap? :)

OpenStudy (anonymous):

all right let me try.

OpenStudy (gorica):

may I write first and then you correct my mistakes?

OpenStudy (anonymous):

fine

OpenStudy (gorica):

\[e_n=-\sin \theta i+\cos \theta j\] \[e_t=\cos \theta i +\sin \theta j\] \[e_n '=-\theta '\cos \theta i-\theta '\sin \theta j=-\theta'e_t\] \[r=-lcos \theta e_n\] \[r'=l \theta'\sin \theta e_n-lcos \theta e_n '=l \theta'\sin \theta e_n+l \theta 'e_t\] If this is good, that, by what my professor wrote, \[l \theta '\sin \theta e_n\] is supposed to be zero.

OpenStudy (souvik):

you guys want this?\(\huge\vec{f}\)

OpenStudy (souvik):

i mean the vector sign over a letter...!

OpenStudy (gorica):

yes, and dot too.

OpenStudy (souvik):

i know only the vector sign..\(\vec{f}\) slash-first bracket open-slash-"vec"-second bracket open-"f"-second bracket close-slash-first bracket close slash means "\" hope u understand....!

OpenStudy (souvik):

for more details in latex.... http://openstudy.com/users/souvik#/updates/51d58ebbe4b059f463fe7608

OpenStudy (anonymous):

It is all too confusing! What is 'l' you are using above?

OpenStudy (gorica):

sorry :D I wrote l as I did on paper... it's length of a string. I was supposed to write s.

OpenStudy (souvik):

u can do this...left click on the f vector i posted then "show math as" and then "tex commands "... here u can see what i wrote..

OpenStudy (souvik):

do not write those what i told.. type this... \(\vec{f}\] just write ')' in place of ']'

OpenStudy (souvik):

@gorica

OpenStudy (gorica):

@souvik thanks, I got it :)

OpenStudy (souvik):

:) too difficult to remember those code words...

OpenStudy (vincent-lyon.fr):

What do you mean by "position vector equation" ? Could you give an example using a simpler case?

OpenStudy (anonymous):

This is all i know. Position vector ,r is written as r=(distance of point from origin)*unit vector pointing from origin to the point. here distance is 's' and unit vector u are using is \(\epsilon_{t}\) So, \(r=s\epsilon_{t}\) The normal vector, \(\epsilon_{n}\) is perpendicular to \(\epsilon_{t}\) and related as follows (u proved somewhat similar above) \(\epsilon_{t}'\)=\(\theta '\)\(\epsilon_{n}\) and \(\epsilon_{n}'\)=-\(\theta '\)\(\epsilon_{t}\) So \(r'=s\theta' \epsilon_{n}\) Similarly \(r''=s\theta''\epsilon_{n} -s\theta'^{2}\epsilon_{t}\) Sorry, i cannot do more!

OpenStudy (vincent-lyon.fr):

Is it its equation wrt time? Is it the (differential) equation of motion?

OpenStudy (gorica):

@Diwakar thank you, that's it. Now I see my mistake, I used a projection on y-axes... Thanks!